A block (mass 5 kg) oscillates on a spring (spring constant 180 N/ m). At one moment, the block is 10 cm from its equilibrium position and is moving with a speed of 80 cm/s away from the equilibrium position. Determine the amplitude of this oscillation.
Answer #1

Let the amplitude of the oscillation is A.

Let the period of the oscillation is T.

Given that the mass of the block m=5 kg.

The spring constant k=180 N/m.

The speed of the block v=80 cm/s=0.80 m/s.

The position of the block x=10 cm=0.10 m.

Now we know the period of the oscillation

T=2\pi \sqrt{\frac{m}{k}}

Now we know the speed of the block at any point is

  v=\pm \omega \sqrt{A^{2}-x^{2}}

  or\: \: \: \: \: \: v= \omega \sqrt{A^{2}-x^{2}}

or\: \: \: \: \: \: v=\frac{2\pi}{T} \times \sqrt{A^{2}-x^{2}}

or\: \: \: \: \: \: v=\frac{2\pi}{2\pi \sqrt{\frac{m}{k}}} \times \sqrt{A^{2}-x^{2}}

  or\: \: \: \: \: \: v=\sqrt{\frac{k}{m}} \times \sqrt{A^{2}-x^{2}}

     or\: \: \: \: \: \: v^{2}=\frac{k}{m}\times\left ( A^{2}-x^{2} \right )

  or\: \: \: \: \: \: A^{2}-x^{2} =\frac{mv^{2}}{k}

  or\: \: \: \: \: \: A^{2} =\frac{mv^{2}}{k}+x^{2}

     or\: \: \: \: \: \: A =\sqrt{\frac{mv^{2}}{k}+x^{2}}

or\: \: \: \: \: \: A =\sqrt{\frac{5\times (0.80)^{2}}{180}+0.10^{2}}

\therefore \: \: \: \: \: \: A =0.1666666667\approx 0.167\: \: \: \: m

\therefore \: \: \: \: \: \: A =16.7\: \: \: \: cm

\therefore The amplitude of the oscillation is 0.167 m or 16.7 cm approx.